Variable Cost Concept of Product Pricing Voice Com, Inc., produces and sells cellular phones. The costs of producing and selling 8,000 units of cellular phones are as follows: Variable costs: Fixed costs: Direct materials $ 65 per unit Factory overhead $295,800 Direct labor 30 Selling and admin. exp. 103,950 Factory overhead 20 Selling and admin. exp. 15 Total $130 per unit Voice Com desires a profit equal to a 15% rate of return on invested assets of $455,000. Assume that Voice Com, Inc., uses the variable cost concept of applying the cost-plus approach to product pricing. a. Determine the variable costs and the variable cost amount per unit for the production and sale of 8,000 units of cellular phones. Total variable costs $ 1,040,000 Variable cost amount per unit $ 130 b. Determine the variable cost markup percentage for cellular phones. % c. Determine the selling price of cellular phones. Round to the nearest cent. $ per phone

I am struggling with number B and C.
Answer #1

Answer:

b.

Desired Profit is 15% of $455,000= $68,250

Total Fixed Cost = $ 295,800 + $ 103,950 = $ 399,750

Assume 'x' is the selling price then sales amount for 8,000 units is 8,000 x

                                 (Fixed Cost + Desired Profit) /Contribution=Required Sales

                                   ($ 399,750 + $68,250)/(x-130)=8000

                                                                 $468,000=8000x-130*8000

                                                                   $468,000+1,040,000=8000x

                                                                    x=1,508,000/8,000=188.5

Variable cost mark up %

                         Selling Price 188.5 100%

            Less :Variable cost    130          66.67%

                                           ______

Contribution per unit              58.5 33.33%

c. Selling Price - $188.5